# On the Relationship between Space-Time Accessibility and Leisure Activity Participation

### v1

@@ 3.3 Socio-economic attributes
The Individuals dataset provides a rich set of sociodemographic, household, and mobility-related attributes. It includes basic identifiers and residence information such as municipality codes, as well as demographic variables like sex, age, education level, and socio-professional category. We also augment the attributes with the poverty rate at the IRIS (9-digit) zone level, used as a proxy for income in 2021. This measure corresponds to the share of individuals living in households whose standard of living—after accounting for taxes, transfers, and social benefits—falls below 60% of the national median disposable income (Institut national de la statistique et des études économiques (2024), INSEE). Household composition is captured through the number of persons and their age distribution, along with housing type. Mobility resources and constraints are represented by indicators of driving licence ownership and car availability, complemented by information on access to two-wheelers, bicycles, e-scooters, and other mobility devices. The dataset also records subscriptions to public transport and other mobility services, alongside a statistical weighting coefficient for representativeness.
@@ 3.5 Point of interest (POI) data
POI data were collected via the Overture API, restricted to the geographical extent of the trip records and filtered for a confidence level above 0.7 (Overture Maps Foundation, 2023). We then labeled these POIs into two categories aligned with the activity purposes recorded in the trip data, Social & Leisure (67 k), consolidated through a combination of GPT-4o–assisted classification and manual validation.
@@ 4.1 Space-time accessibility
In this study, we operationalize individual-based capability set as space–time accessibility (SPA). Specifically, we focus on the cardinal accessibility dimension, which accounts for the number of opportunities (leisure POIs, including parks, etc.) that individuals can feasibly reach within their available time budget (Saraiva and Barros, 2022).
Figure 3: Space–time accessibility (SPA). Adapted from Saraiva and Barros (2022).
The measure builds on the space-time feasibility concept (Kwan, 1999) and is modified based on the cardinal individual-based accessibility measure proposed in (Saraiva and Barros, 2022). Each individual is assumed to have a fixed time budget $tb_{i}$, defined as the residual time available for discretionary activities after completing mandatory activities (e.g., at home and work). An activity $k$ is considered accessible to individual $i$ as if he/she does a one-stop trip chaining in between their trip from work to home, as more commonly found in the population (McGuckin and Murakami, 1999). In other words, the total time required to (i) travel from home to work ($t_{hw}$), (ii) travel from work to $k$ ($t_{wk}$), and (iii) return from $k$ to home ($t_{kh}$) does not exceed the time budget.
@@
  • $tb_{i}$ is the time budget of individual $i$.
The overall space–time accessibility $A_{i}^{k}$ of individual $i$ is then given by the sum of all feasible opportunities under the $k$ category:
$A_{i}^{k}=\sum a^{k}_{i}$ (2)
In this study, we define the individual travel time budget $tb_{i}$ as 90 minutes for all individuals. Commuting time ($t_{hw}$) is empirically estimated using the weighted median of observed commuting trips. Discretionary activity $k$ is defined as a visit to leisure-related locations (e.g., bars).
For calculating $t_{wk}$ and $t_{kh}$, we fix the departure time at 17:00, resembling the work-to-home trip chaining. Travel times between home, work, and candidate leisure POIs were computed for both public transit and car, using GTFS schedules and road network data, implemented through the r5r package (Pereira et al., 2021).
@@
While the space-time accessibility construct remains somewhat opaque, realistic in formulation, the actual spatial choices in leisure activity participation can differ. To gain a finer-grained understanding of how accessibility structures behavior, we assess whether individuals tend to select activity locations that align with what their modeled opportunity landscape (SPA) predicts as desirable, or whether actual behavior departs from this logic.
This behavioral analysis focuses on the alignment between the ranked set of feasible locations SPAi (adding up to $A_{i}$) and the actual visited locations for each individual $i$. By comparing observed choices against random draws from SPAi, we test for selectivity—that is, whether individuals systematically prefer higher-ranked locations.
We implement a rank-based comparison. We evaluate whether individuals tend to choose better-ranked locations (e.g., those that take a shorter time to reach) than would be expected by chance. For each individual $i$, we construct their SPAi—a ranked list of feasible locations (H3 hexagons at resolution 8, $\sim$0.74 km2) sorted from 1 (best) to $N_{i}$ (worst) based on the modeled opportunity score (e.g., time left in the travel time budget $tb_{i}-t_{hw}-t_{wk}-t_{kh}$).
We then compare the actual visited locations $V_{i}$ to this ranking. First, we log the share of locations in $V_{i}$ that fall outside SPAii. Second, the test statistic is the average rank of these visited locations fall within SPAii:
$T_{\text{mean},i}^{\text{act}}=\frac{1}{k_{i}}\sum_{\ell\in V_{i}}\text{rank}(\ell)$ (3)
To generate a null distribution, we perform $B=1000$ random draws of $k_{i}$ locations from SPAii, computing $T_{\text{mean},i}^{(b)}$ for each. We then derive both the empirical $p$-value:
$p_{i}=\frac{1+\#\left\{b:T_{\text{mean},i}^{(b)}\leq T_{\text{mean},i}^{\text{act}}\right\}}{B+1}$ (4)
@@
$d_{i}=\frac{T_{\text{mean},i}^{\text{act}}-\mu_{i}^{\text{null}}}{\sigma_{i}^{\text{null}}}$ (5)
@@
4.3 Mobility quantification
We quantify individual mobility along two dimensions: trip making, captured by daily total travel time, and third-place activity participation, characterised by the location diversity. Because the travel diary lacks detailed labeling of third-place activities, we approximate third-place participation using all leisure activities (’LEISURE’). We compute a diversity measure based on the distribution of leisure activity locations. For each individual $i$, we count the number of visits to each distinct location. Let $n$ denote the total number of leisure visits, and let $K$ denote the number of distinct locations visited. Defining $p_{k}=n_{k}/n$ as the relative frequency of visits to location $k$ ($k=1,\dots,K$), we compute the Hill number of order $q=1$, which is given by:
@@ 4.3 Mobility quantification
which reflects the effective number of equally frequent locations visited, and accounts for both richness and evenness in the individual’s leisure activity location distribution.
@@ 4.4 Structural equation modeling specification
Based on the data summarized in Table 1, we estimate a structural equation model (SEM) linking (i) individual attributes, including one lifestyle metric on active transport mode use, (ii) transport mode, (iii) space-time accessibility (SPA), (iv) trip making, and (v) leisure activity participation. Let $i$ index individuals. The observed variables are:
@@
  • $H_{1i}$: leisure activity participation, measured by the Hill number of order $q=1$.
Table 1: Descriptive statistics of the 2,415 commuting individuals. Values represent weighted means (continuous variables) and weighted percentages (categorical variables).
@@
Before estimating the structural equation model, it is important to specify a directed acyclic graph (DAG) that encodes the theoretical and causal assumptions about the relationships among the variables. The DAG serves as a transparent foundation for identifying potential confounding paths, testing conditional independence implications, and ensuring that the SEM structure is both theoretically grounded and empirically justified.
After testing for testable implications, the final directed acyclic graph (DAG) reflects the theoretically informed and empirically tested structure of the relationships among individual attributes, transport mode, capability set, trip making, and activity participation (Figure 4). Individual characteristics are modeled as exogenous factors that shape both transport mode choice and accessibility (capability set), as well as exerting direct influences on trip-making and activity participation. Transport mode mediates part of these effects by influencing accessibility, trip-making, and participation directly. The capability set, representing the feasible space–time leisure opportunities available to individuals, is affected by both personal attributes and transport mode, and in turn drives patterns of trip-making and participation. Trip-making operates as an intermediate behavioral mechanism linking accessibility to leisure activity participation. Finally, leisure activity participation is conceptualized as the outcome of multiple interacting pathways: directly from individuals, from transport behavior, and indirectly via both the capability set and trip-making. This structure balances theoretical expectations from time-geography and accessibility theory with empirical evidence from conditional independence testing, ensuring that only statistically supported pathways are retained.
Individualattributes($\mathbf{Z}_{i}$)Transportmode($\mathbf{M}_{i}$)Capability Set($A_{i}$)Trip making($B_{i}$)Activityparticipation($H_{1i}$)
Figure 4: Pathway structure. This directed acyclic graph shows the hypothesized and empirically validated pathways. Exposure=Capability set, Outcome=Activity participation, Black arrows=Causal paths, Brown arrows=Biasing paths.
In SEM, the structural relations in Figure 4 are specified as follows:
@@
The model captures both direct associations of SPA with leisure participation ($c$), and indirect associations transmitted through travel behavior pathways ($a_{tt}b_{tt}$). In addition, sociodemographic attributes $\mathbf{Z}_{i}$ contribute to leisure activity location diversity both directly (coefficients $\mathbf{d}$) and indirectly via transport mode ($\mathbf{M}_{i}$), SPA ($A_{i}$), and trip making ($B_{i}$).
@@ 5.1 Descriptive statistics
Figure 5 presents a spatial overview of space–time accessibility (SPA), trip making, and leisure activity participation across the study region, highlighting the geographic variations. Figure 5a shows the proportion of inhabitants with SPA >0>0, indicating the share of individuals who have feasible opportunities beyond their fixed home and work locations. Higher values are concentrated in the northern central urban areas and scattered across select suburban zones, suggesting greater access to leisure activity opportunities in both dense and well-connected peripheral locations, mainly by car. Figure 5b maps the average total travel time, with higher values concentrated in peripheral areas. This pattern underscores the role of trip-making as a necessity in facilitating activity demand rather than a choice, driven by limited service availability on the outskirts. Figure 5c visualizes leisure activity location diversity. Compared to SPA and total travel time, diversity values show weaker spatial clustering, as they are more strongly shaped by individual-level differences in activity behavior. Together, the maps suggest that while core urban areas tend to offer greater capability sets and diverse leisure opportunities, spatial disparities in accessibility, mobility, and activity participation persist across individuals.
@@
Figure 5: Space–time accessibility (SPA) and mobility metrics on the map. a, Share of inhabitants with SPA > 0. b, Total travel time (daily, min). c, Leisure activity diversity (weighted median of each zone’s inhabitants).
@@
We further examine how SPA relates to total travel time and leisure activity participation, stratified by main transport mode (car vs. public transit) in Figure 6. Figure 6a shows that individuals with SPA $\neq 0$ have substantially lower travel times, suggesting that feasible opportunities beyond home and work enable more efficient activity engagement. Figure 6b illustrates a clear negative association between SPA and travel time for both car and public transit users, with the effect more pronounced among car users. Figure 6c shows that leisure activity diversity increases with accessibility, particularly for public transit users. Taken together, these results indicate that higher accessibility not only facilitates efficiency but also broadens the diversity of leisure participation, though the magnitude and pathways of these effects differ between car and public transit users.
@@
The correlational results (Figure 6) support our assumptions regarding the interdependence of individual attributes, transport mode choice, and space–time accessibility (Figure 2). They further suggest that an individual’s space–time accessibility shapes both trip-making behavior and actual leisure activity participation. SEM outcomes further disentangle their relationships.
@@
Figure 6: Space–time accessibility $A_{i}$ and trip making and activity behaviors across transport modes. a, Total travel time and leisure activity diversity, comparing those with $A_{i}$ = 0 (grey) and $A_{i}$ > 0 (green). b, Total travel time as a function of $A_{i}$ for car and public transit users ($A_{i}$ > 0). c, Leisure activity diversity ($A_{i}$ > 0) by mode. Error bars indicate bootstrap median estimation errors.

### v2

@@ 3.3 Socio-economic attributes
The data from the EMG 2023 survey provides a rich set of sociodemographic, household, and mobility-related attributes. It includes basic identifiers and residence information such as municipality codes, as well as demographic variables like sex, age, education level, and socio-professional category. We also augment the attributes with the poverty rate at the IRIS (9-digit) zone level, used as a proxy for income in 2021. This measure corresponds to the share of individuals living in households whose standard of living—after accounting for taxes, transfers, and social benefits—falls below 60% of the national median disposable income (INSEE_RevenusPauvreteIris2021). Household composition is captured through the number of persons and their age distribution, as well as the type of housing. Mobility resources and constraints are represented by indicators of driving licence ownership and car availability, complemented by information on access to two-wheelers, bicycles, e-scooters, and other mobility devices. The dataset also records subscriptions to public transport and other mobility services, alongside a statistical weighting coefficient for representativeness.
@@ 3.5 Point of interest (POI) data
POI data were collected via the Overture API, restricted to the geographical extent of the trip records, and filtered for a confidence level above 0.7 (overturemaps). We then extracted POIs and their primary and secondary labels that are aligned with the activity purposes recorded in the trip data, Social & Leisure (67 k), consolidated through a combination of GPT-4o assisted classification (Essential needs, Health services, Education, Civic and utility, Social & Leisure, and Other) and manual validation.
@@ 4 Methodology
In this section, we outline the key components of the four methods used in our analysis. We begin by describing the calculation of space–time accessibility (Section 4.1), followed by the approaches used to assess how well individuals’ capability sets (reflected in their space–time accessibility) align with their observed spatial choices of leisure destinations (Section 4.2). We then present the methods for quantifying mobility through total travel time and participation in third-place activities (Section 4.3). Finally, we introduce the structural equation model and its specification to disentangle the complex relationship between space-time accessibility and leisure activity participation (Section 4.4). The source code of this study is available on GitHub.
@@ 4.1 Space-time accessibility
In this study, we operationalize individual-based capability set as space–time accessibility (STA). Specifically, we focus on the cardinal accessibility dimension, which accounts for the number of opportunities (leisure POIs, including parks, etc.) that individuals can feasibly reach within their available time budget (saraiva2022accessibility). In other words, it counts the number of opportunities located within a person’s Potential Path Area (PPA), which is the geographical area a person can reach given her space-time constraints (saraiva2022accessibility).
Figure 3: Space–Time Accessibility (STA). The light-green region represents the Potential Path Area (PPA), i.e., all locations reachable given the individual’s time budget and travel constraints between home and work. Opportunities within this area (dark green) are accessible, while those outside it (light green) are not. The direct home–work path is shown in bold orange, and a feasible work–activity–home path is shown in blue. This illustrates the STA of an individual, defined as the set of feasible opportunities that fall within the individual’s PPA. Adapted from saraiva2022accessibility.
The measure builds on the space-time feasibility concept (kwan1999gender) and is modified based on the cardinal individual-based accessibility measure proposed in (saraiva2022accessibility). Each individual is assumed to have a fixed total travel time budget $tb_{i}$ and use their revealed main transport mode (car or public transport). An activity destination $k$ is considered accessible to individual $i$ if the time required to visit $k$ does not exceed the individual’s time budget. In other words, the total time required to travel from home to work ($t_{hw}$), from work to $k$ ($t_{wk}$), and return from $k$ to home ($t_{kh}$) does not exceed the time budget, as if he/she does a one-stop trip chaining in between their trip from work to home, as more commonly found in the population (mcguckin1999examining). In other words, the total time required to (i) travel from home to work ($t_{hw}$), (ii) travel from work to $k$ ($t_{wk}$), and (iii) return from $k$ to home ($t_{kh}$) does not exceed the time budget.
@@
  • $tb_{i}$ is the travel time budget of individual $i$.
With the overall space–time accessibility (STAi, the capability set), the number of opportunities within STAi is $A_{i}$ for individual $i$, given by the sum of all feasible opportunities under certain category:
$A_{i}=\sum a^{k}_{i}$ (2)
In this study, we set each individual’s travel time budget $tb_{i}$ to 90 minutes, based on the empirical median of 82 minutes observed in our dataset. Commuting time ($t_{hw}$), travel time from home to work, is empirically estimated using the weighted median of observed commuting trips. Discretionary activity $k$ is defined as a visit to leisure-related locations (e.g., bars).
For calculating $t_{wk}$ and $t_{kh}$, we fix the departure time at 17:00, which resembles the work-to-home trip chaining scenario with an added leisure activity stop on the way from work to home. Travel times between home, work, and candidate leisure POIs were computed for both public transit and car, using GTFS schedules and road network data, implemented through the r5r package r5r.
@@
To gain a finer-grained understanding of how accessibility shapes behavior, we assess whether individuals tend to select activity locations that align with what their modeled opportunity landscape (STA) predicts as desirable, or whether actual behavior departs from this logic.
This behavioral analysis focuses on the alignment between the ranked set of feasible locations STAi (adding up to $A_{i}$) within individuals’ potential path area and the locations they actually visited (adding up to $K_{i}$). By comparing observed choices against random draws from STAi, we test for selectivity—that is, whether individuals systematically prefer most accessible locations (see Figure 4).
Figure 4: Space–Time Accessibility (STA) set vs. visited leisure locations. One example individual. Home and work icons show where the person lives and works. Hexagons are at resolution 8, $\sim$0.74 km2.
We evaluate whether individuals tend to visit the activities that are the most accessible (e.g., those that take a shorter time to reach) using a ranked comparison: For each individual $i$, we operationalize their STAi as a ranked list of feasible locations (H3 hexagons at resolution 8, $\sim$0.74 km2). We focus on H3 hexagons rather than raw GPS coordinates for two reasons. First, not all locations labeled as leisure in the GPS data correspond to entries in the leisure POI dataset extracted from Overture Maps. Second, the GPS coordinates have been preprocessed by the data provider to H3 level-10 centroids as a privacy measure, making precise point-level POI matching infeasible. Consequently, POI assignment must rely on an appropriate spatial unit rather than exact-coordinate matching. These spatial units (H3 hexagons at resolution 8) are further sorted from 1 (best = most accessible, lowest travel time increase) to $N_{i}$ (worst = least accessible, highest travel time increase).
We then compare the actual visited locations $V_{i}$ to this ranking. First, we calculate the share of locations in $V_{i}$ that fall outside STAi. Second, we define the test statistic as the average rank of the visited locations falling within STAi:
$T_{\text{mean},i}=\frac{1}{K_{i}}\sum_{\ell\in V_{i}}\text{rank}(\ell)$ (3)
To generate a null distribution, we perform $B=1000$ random draws of $K_{i}$ locations from STAi, computing $T_{\text{mean},i}^{(b)}$ for each. We then derive both the empirical $p$-value:
$p_{i}=\frac{1+\#\left\{b:T_{\text{mean},i}^{(b)}\leq T_{\text{mean},i}\right\}}{B+1}$ (4)
@@
$d_{i}=\frac{T_{\text{mean},i}-\mu_{i}^{\text{null}}}{\sigma_{i}^{\text{null}}}$ (5)
@@
4.3 Mobility and activity quantification
We consider two dimensions: total travel time, and third-place activity participation, captured by the diversity of locations where leisure activities occur. Because the travel diary lacks detailed labeling of third-place activities, we approximate third-place participation using all leisure activities (’LEISURE’). For each individual $i$, we count the number of visits to each distinct leisure location (H3 hexagons at resolution 10) during their data collection period. Let $n$ denote the total number of leisure visits, and let $K$ denote the number of distinct leisure locations visited. Defining $p_{k}=n_{k}/n$ as the relative frequency of visits to leisure location $k$ ($k=1,\dots,K$), we compute the Hill number of order $q=1$, which is given by:
@@ 4.3 Mobility and activity quantification
which reflects the effective number of equally frequent locations visited, and accounts for both richness and evenness in the individual’s leisure location distribution.
@@ 4.4 Structural equation modeling specification
Based on the data summarized in Table 1, we estimate a structural equation model (SEM) linking (i) individual attributes, (ii) transport mode, (iii) space-time accessibility value $A_{i}$, (iv) total travel time, and (v) leisure location diversity. Let $i$ index individuals. The observed variables are:
  • i.
@@
  • v.
    $H_{1i}$: leisure activity participation, measured by the Hill number of order $q=1$, quantifying the diversity of leisure locations visited.
Table 1: Descriptive statistics of the 2,415 commuting individuals. Values represent weighted means (continuous variables) and weighted percentages (categorical variables). $A_{i}$ is log-transformed for positive values, while zeros are retained.
@@
Before estimating the structural equation model, it is important to specify a directed acyclic graph (DAG) that encodes the theoretical and causal assumptions about the relationships among the variables. The DAG serves as a transparent foundation for identifying potential confounding paths, making the assumptions of the model explicit, and ensuring that the SEM structure is both theoretically grounded and empirically justified (huntington2021effect). We iteratively evaluated which relationships could be pruned by testing the key conditional independence implications suggested by our theoretical framework. We assessed whether certain connections between variables—such as those linking individual attributes, transport mode, capability set, mobility, and activity participation—could be removed without contradicting the observed data. This process allowed us to identify which dependencies were empirically warranted and which were not, ensuring that the resulting DAG retained only the pathways supported by theory and evidence.
The final directed acyclic graph (DAG) reflects the theoretically informed and empirically tested structure of the relationships among individual attributes, transport mode, capability set, mobility, and activity participation (Figure 5). Individual characteristics are modeled as exogenous factors that shape both transport mode choice and accessibility (capability set), as well as exerting direct influences on trip-making and activity participation. Transport mode mediates part of these effects by directly influencing accessibility, trip-making, and participation. The capability set, representing the feasible space–time leisure opportunities available to individuals, is affected by both personal attributes and transport mode, and in turn drives patterns of trip-making and participation. Trip-making operates as an intermediate behavioral mechanism linking accessibility to leisure activity participation. Finally, leisure activity participation is conceptualized as the outcome of multiple interacting pathways: directly from individuals, from transport behavior, and indirectly via both the capability set and trip-making. This structure balances theoretical expectations from time-geography and accessibility theory with empirical evidence from conditional independence testing, ensuring that only statistically supported pathways are retained.
Individualattributes($\mathbf{Z}_{i}$)Transportmode($\mathbf{M}_{i}$)STA value($A_{i}$)Mobility($B_{i}$)Activityparticipation($H_{1i}$)
Figure 5: Pathway structure. This directed acyclic graph shows the hypothesized and empirically validated pathways. Exposure=STA value, Outcome=Activity participation, Black arrows=Causal paths, Brown arrows=Biasing paths.
In SEM, the structural relationships in Figure 5 are specified as follows:
@@
The model captures both direct associations of STA with leisure participation ($c$), and indirect associations transmitted through mobility pathways ($a_{tt}b_{tt}$). In addition, sociodemographic attributes $\mathbf{Z}_{i}$ contribute to leisure location diversity both directly (coefficients $\mathbf{d}$) and indirectly via transport mode ($\mathbf{M}_{i}$), STA value ($A_{i}$), and total travel time ($B_{i}$).
@@ 5 Results and discussion
In this section, we first present descriptive statistics and visualizations of the key variables, including the relationships between space–time accessibility and mobility or activity behaviors (Section 5.1). We then report the results assessing the alignment between the modeled capability set and individuals’ actual spatial choices of leisure destinations (Section 5.2). Finally, we present the modeling outcomes (Section 5.3), detailing the quantified relationships and how different components link to space–time accessibility and shape leisure location diversity.
@@ 5.1 Descriptive statistics
Figure 6 presents a spatial overview of space–time accessibility value $A_{i}$ (log-transformed), total travel time, and diversity of visited leisure locations across the study region, highlighting the geographic variations. Figure 6a shows the proportion of inhabitants with STA >0>0, indicating the share of individuals who have feasible opportunities beyond their fixed home and work locations. Higher values are concentrated in the northern central urban areas and scattered across select suburban zones, suggesting greater access to leisure activity opportunities in both dense and well-connected peripheral locations, mainly by car. Figure 6b maps the average total travel time, with higher values concentrated in peripheral areas. This pattern underscores the role of mobility as a necessity in facilitating activity demand rather than a choice, driven by limited service availability on the outskirts. Figure 6c visualizes leisure location diversity. Compared to STA and total travel time, diversity values show weaker spatial clustering, as they appear to be more strongly shaped by individual-level differences in activity behavior.
@@
Figure 6: Space–time accessibility value and mobility metrics on the map. a, Share of inhabitants with AA > 0. Gray areas indicate zero percent. b, Total travel time (daily, min). c, Leisure location diversity (weighted median of each zone’s inhabitants).
@@
We further visualize how STA relates to total travel time and leisure activity participation, stratified by main transport mode (car vs. public transit) in Figure 7. Figure 7a shows that individuals with $A>0$ have substantially lower travel times, suggesting that feasible opportunities beyond home and work link to more efficient activity engagement, i.e., people travel less time to meet their activity demand. Figure 7b illustrates a modest but statistically significant negative association between STA and travel time for both car (Spearman correlation coefficient r=0.18r=-0.18, $p=0.01$) and public transit users (r=0.13r=-0.13, $p=0.04$), with the effect more pronounced among car users. As shown in Figure 7c, leisure location diversity is positively associated with accessibility among public transit users ($r=0.13$, $p=0.05$), whereas no such relationship is observed for car users.
@@
The correlational results (Figure 7) support our assumptions regarding the interdependence of individual attributes, transport mode choice, and space–time accessibility value (Figure 2). They further suggest that an individual’s space–time accessibility shapes both mobility behavior and actual leisure activity participation. SEM outcomes further disentangle their relationships.
@@
Figure 7: Space–time accessibility value $A_{i}$ and mobility and activity behaviors across transport modes. a, Total travel time (top) and leisure location diversity (bottom), comparing those with $A_{i}$ = 0 (grey) and $A_{i}$ > 0 (green). b, Total travel time as a function of $A_{i}$ for car and public transit users ($A_{i}$ > 0). c, Leisure location diversity ($A_{i}$ > 0) by mode. Error bars indicate bootstrap median estimation errors.